OpenStudy (anonymous):

please help me!!! let log(D^2/C^3*A) when logA=3 logB=2 logD=5

OpenStudy (campbell_st):

well the law for division in logs \[\log(\frac{a}{b}) = \log(a) - \log(b)\] and the law for multiplication is \[\log(a \times b) = \log(a) + \log(b)\] and lastly the law for powers \[\log(a^b) = b \times \log(a)\] just apply the laws to simplify the question then substitute the given values.

OpenStudy (anonymous):

still did not get an answer please help me through the steps

OpenStudy (campbell_st):

ok.. so its \[\log(D^2) - \log(C^3) + \log(A)\] so just apply the power law... and then substitute your values.

OpenStudy (anonymous):

i got .97197 but the four answers they give you are 0.549, 4.44,1, and 19 which one would it be?

OpenStudy (campbell_st):

well its interesting how you got your answer as there is no information about logC

OpenStudy (anonymous):

i that is why i am confused